OpenStudy (anonymous):

If f(x) = x-5/6 and g(x) = 6x + 5, find (g,f)(x) a. x b. x + 10 c. x - d. 6x + 25

OpenStudy (anonymous):

(g,f)(x) = g(f(x)) ?

OpenStudy (anonymous):

TYPO::: c. x-5/6

OpenStudy (anonymous):

These are inverses of each other so the answer is A.

OpenStudy (anonymous):

yea.. i got messed up on this one

OpenStudy (eyust707):

g(f(x)) means plug f(x) into g

OpenStudy (eyust707):

so you see all the x's in g? change those to ( x- 5/6)

OpenStudy (eyust707):

6(x - 5/6) + 5 = g(f(x)) ... i think lol (6x - 5) + 5 = g (f(x)) 6x = g(f(x))

OpenStudy (eyust707):

not 100 percent sure tho

OpenStudy (anonymous):

it is x. :)

OpenStudy (anonymous):

You should just get X because they are inverses of each other.

OpenStudy (eyust707):

you would if it was 1/6 x

OpenStudy (eyust707):

ohhh hehe the whole thing is divided by 6

OpenStudy (anonymous):

yeah

OpenStudy (eyust707):

shoulda been ( ) in there =P

OpenStudy (anonymous):

thanks all :)